# Recovering contact forms from boundary data

### v5

@@ 5. On contact forms and contact vector fields on manifolds with boundary
Let $\alpha$ be a differential $k$-form on a smooth manifold $X$ and let $Z\subset X$ be a smooth submanifold. We use somewhat unconventional notations: we write “$\alpha|_{Z}$” for the restriction of the section $\alpha:X\to\bigwedge^{k}T^{\ast}X$ to the section $\alpha|:Z\to\bigwedge^{k}T^{\ast}Z$, and we write “$\alpha||_{Z}$” for the restriction of $\alpha$ to the section of the bundle $\bigwedge^{k}T^{\ast}X|_{Z}\to Z$.
@@
A smooth vector field $u$ on $X$ is called contact field for a given contact $1$-form $\beta$ if $\mathcal{L}_{u}\beta=\lambda\cdot\beta$ for some smooth function $\lambda:X\to\mathbb{R}$. $\diamondsuit$
@@
To derive Corollary 5.2 of Theorem 4.1 we will rely on a couple of lemmas.
@@
Using that $\mathcal{L}_{v_{\beta}}\beta\equiv 0$ and $v_{\beta}\,\rfloor\,\beta=1$, this is equivalent to the condition $1\cdot dh+\mathcal{L}_{w}\beta=\lambda\cdot\beta,$ or, by Cartan’s formula, to the condition
@@
By Lemma 5.3, $u=h\cdot v_{\beta}+w$ is also tangent to $\partial_{1}X$, since $h|_{\partial_{1}X}=0$. Thus, we getting the following claim.
@@
$\bullet$ Let $X=\{x^{2}+y^{2}+z^{2}\leq 1\}\subset\mathbb{R}^{3}$ and $\beta=dz+x\,dy$. The Reeb vector field $v_{\beta}$ is $\partial_{z}$. It is tangent to the $2$-sphere $\partial_{1}X$ along its equator $E=\{z=0\}\cap X$. Thus, the restriction of $\beta$ to $E$ is equal to the restriction of the form $x\,dy$ to $E$. If $\theta$ is the natural parameter on the circle $E$, then $\beta|_{E}=cos^{2}(\theta)\,d\theta$. It is non-negative everywhere on $E$ and vanishes at $\theta=\pm\pi/2$, or at $(x,y,z)=(0,\pm 1,0)$.
@@
The $u$-flow preserves the distribution $\xi_{\beta}$ and the sphere $\partial_{1}X$. It has exactly at two fixed points $(0,0,\pm 1)$.
$\bullet$ We pick the shell $\{1\leq x^{2}+y^{2}+z^{2}\leq 4\}$ for the role of $X$ and the same contact form $\beta=dz+x\,dy$. Then $\partial_{2}^{+}X(v_{\beta})$ is the equator $E_{1}=\{x^{2}+y^{2}+z^{2}=1\}\cap\{z=0\}$ and $\partial_{2}^{-}X(v_{\beta})$ is the equator $E_{2}=\{x^{2}+y^{2}+z^{2}=4\}\cap\{z=0\}$. The same computation leads to $\beta|_{E_{1}}\geq 0$ and $\beta|_{E_{2}}\leq 0$. In other words, ±$\pm\beta|_{\partial_{2}^{\pm}X(v_{\beta})}\geq 0$.
Now the function $h=(x^{2}+y^{2}+z^{2}-4)(x^{2}+y^{2}+z^{2}-1)$. We leave to the reader to solve the equation $w\,\rfloor\,d\beta=-dh+dh(v_{\beta})\cdot\beta$ for $w\in\xi_{\beta}$.
@@
[…]
@@
  • For an even $j$,  $\beta|_{\mathsf{int}(\partial_{j}^{+}X(v_{\beta}))}$ is a contact form away from a zero measure set. Moreover,
@@
  • For an odd $j$, $d\beta|_{\mathsf{int}(\partial_{j}^{+}X(v_{\beta}))}$ is a symplectic form away from a zero measure set. Moreover,
    ±$\pm(d\beta)^{\frac{2n+1-j}{2}}\big{|}_{\partial_{j}^{\pm}X(v_{\beta})}\geq 0,\text{ and }(d\beta)^{\frac{2n+1-j}{2}}\big{|}\big{|}_{\partial_{j+1}X(v_{\beta})}\equiv 0\;.$
@@
Recall that the Gray Stability Theorem [Gray] claims that, for any smooth family $\{\xi_{t}\}_{t\in[0,1]}$ of contact structures on a closed smooth manifold $X$, there exists a smooth isotopy $\{\phi_{t}\}_{t\in[0,1]}$ such that $(\phi_{t})^{\ast}(\xi_{t})=\xi_{0}$ for all $t\in[0,1]$. In fact, we may assume that $X$ is compact in this claim, provided $\dot{\beta}_{t}||_{\partial_{1}X}=dh_{t}||_{\partial_{1}X}$, where $\dot{\beta}_{t}=_{\mathsf{def}}\frac{d}{dt}\beta_{t}$.
@@
Let $X$ be a compact connected smooth $(2n+1)$-dimensional manifold $X$ with boundary. Let $\{h_{t}:\hat{X}\to\mathbb{R}\}_{t\in[0,1]}$ be a family of smooth functions, each of which has 00 as its regular value, $h_{t}^{-1}(0)=\partial_{1}X$, and $h_{t}^{-1}((-\infty,0])=X$.
Consider a family $\{\xi_{t}\}_{t\in[0,1]}$ of cooriented contact structures on $X$, generated by a smooth family of contact forms $\{\beta_{t}\}_{t\in[0,1]}$, such that the $t$-derivative $\dot{\beta}_{t}=dh_{t}$ along $\partial_{1}X$ [^4] The property $\dot{\beta}_{t}||_{\partial_{1}X}=dh_{t}||_{\partial_{1}X}$ implies that $\beta_{t}|_{\partial_{1}X}$ is constant.
Then there is a smooth $t$-family of diffeomorphism $\{\phi_{t}:X\to X\}_{t\in[0,1]}$, such that $(\phi_{t})^{\ast}(\xi_{t})=\xi_{0}$; in particular, $\phi_{1}$ is a contactomorphism for the pair $\xi_{0},\;\xi_{1}$.
@@
Corollary 5.2.
We adopt the notations and hypotheses of Lemma 5.5: let a $t$-family of contact forms $\{\beta_{t}\}_{t\in[0,1]}$ on $X$ be such that the $t$-derivative $\dot{\beta}_{t}||_{\partial_{1}X}=dh_{t}||_{\partial_{1}X}.$ Let $\{\xi_{t}\}_{t\in[0,1]}$ be the corresponding $t$-family of contact structures.
Assume that the Reeb vector field $v_{0}$ of the contact form $\beta_{0}$, is traversing.
Then the Reeb vector field $v_{1}$ of a contact form $\beta_{1}$, is traversing as well. As a result, any $v_{1}$-trajectory is homeomorphic to a closed interval or to a singleton (in particular, $v_{1}$ has no closed or trapped trajectories).
If, in addition, $v_{0}$ is boundary generic, then the causality maps $C_{v_{1}}$ and $C_{v_{0}}$ are conjugate with the help of the diffeomorphism $\phi_{1}^{\partial}=\phi_{1}|:\partial_{1}X\to\partial_{1}X$ as in Lemma 5.5.
Moreover, the knowledge of the boundary data $\{C_{v_{\beta_{0}}},(f^{\bullet}_{0})^{\partial},\beta_{0}^{\partial}\}$ is sufficient for reconstructing the smooth topological type of $\{X,\xi_{1}\}$, provided that $v_{\beta_{0}}$ satisfies Property $\mathsf{A}$.
Proof.
By Lemma 5.5, for an appropriate diffeomorphism $\phi_{1}:X\to X$, we get $\phi_{1}^{\ast}(\beta_{1})=\lambda_{1}\cdot\beta_{0}$. By arguments as in the proof of Corollary 4.7, we get that $v_{1}=(\phi_{1})_{\ast}(\lambda_{1}^{-1}v_{0})$ is the Reeb vector field for $\beta_{1}$. This implies that, with the help of $\phi_{1}:X\to X$, the smooth dynamics of the $v_{0}$-flow and the $v_{1}$-flow are similar. In particular, if $v_{0}$ is traversing, then so is $v_{1}$. If $v_{0}$ is boundary generic, then so is $v_{1}$. For the same reason, the causality maps $C_{v_{1}}$ and $C_{v_{0}}$, which depend only on the conformal classes of vector fields $v_{0}$ and $v_{1}$, are conjugate with the help of $\phi_{1}^{\partial}$.
The last assertion follows by Combining Theorem 4.1 (or Corollary 4.8) with Lemma 5.5, provided that $v_{\beta_{0}}$ satisfies Property $\mathsf{A}$. ∎
@@ 6. Non-squeezing of $2n$-symplectic volumes, Morse wrinkles of Reeb fields, and holography
Let $v_{\beta}$ be a traversing Reeb vector field for a contact $1$-form $\beta$ on $X$. The Reeb diameter of $(X,\beta)$ is defined by
@@
where $\gamma$ runs over the $v_{\beta}$-trajectories. $\diamondsuit$
@@
Consider an orientation-preserving submersion $\Psi:X\to Y$ of two compact equidimentional manifolds. Let $\beta_{Y}$ be a contact form on $Y$ whose Reeb vector field $v_{Y}$ is traversing. We assume that $\Psi(\partial_{1}X)$ is boundary generic relative to $v_{Y}$.
@@
If $\Psi$ is a covering map (i.e., an immersion that is also surjective), then
@@
Since $\Psi$ is submersion and $\Psi(\partial_{1}X)$ is boundary generic relative to $v_{Y}$, for any $v_{Y}$-trajectory $\gamma$ (which is a closed interval or a singleton) the set $\Psi^{-1}(\gamma)$ is either a disjoint union of finitely many closed intervals $\tilde{\gamma}$, or a finite set. Moreover, $\Psi:\tilde{\gamma}\hookrightarrow\gamma$ is an immersion. Thus, $\int_{\tilde{\gamma}}\beta_{X}=\int_{\tilde{\gamma}}\Psi^{\ast}(\beta_{Y})=\int_{\Psi(\tilde{\gamma})}\beta_{Y}\leq\int_{\gamma}\beta_{Y}$. In the case of a covering $\Psi$, we get $\Psi:\tilde{\gamma}\to\gamma$ is a diffeomorphism. In such a case, $\int_{\tilde{\gamma}}\beta_{X}=\int_{\gamma}\beta_{Y}$. Since any $v_{X}$-trajectory $\tilde{\gamma}$ is obtained from some $v_{Y}$-trajectory $\gamma$ as described above, the claims about the Reeb diameters follow. ∎
@@
The argument is a variation on the theme of Santalo formula [S]. For any $y\in\partial_{1}^{+}X(v_{\beta})$ we consider the $v$-trajectory $\gamma_{y}$ through $y$. The $(-v)$-directed projection $p:X\to\partial_{1}^{+}X(v_{\beta})$ is a submersion, away from the set $p^{-1}(\partial_{2}X(v_{\beta}))$ of zero measure in $\partial_{1}^{+}X(v_{\beta})$. Thus, we can use the Dieudonne generalization of the Fubini formula [D]:
@@
together with Definition 6.1, to validate (6.2). By Stokes’ theorem, applied to $\partial_{1}^{+}X(v_{\beta})$, (6.3) follows.
Although this argument is simple, it is somewhat surprising to realize the $(d\beta)^{n}$-volume of the trajectory space $\mathcal{T}(v_{\beta})$ depends only on the form $\beta\wedge(\partial\beta)^{n-1}|_{\partial_{2}X(v_{\beta})}$ on the “equator” $\partial_{2}X(v_{\beta})$. ∎
@@
By (6.3), $\mathsf{av}_{\mathcal{R}}(\beta)\;\leq\;\mathsf{diam}_{\mathcal{R}}(\beta).$diamondsuit
Example 7.1. Let $\beta=dz+x\,dy$ and $X=\big{\{}\frac{x^{2}}{(A/2)^{2}}+\frac{y^{2}}{(B/2)^{2}}+\frac{z^{2}}{(C/2)^{2}}\leq 1\big{\}}$. Then the inequality (6.2) can be written as
@@
For any $\eta\in C^{\infty}(\mathcal{T}(v_{\beta}))$, the contact form $\beta_{1}=\beta+d\eta$ shares with $\beta$ the Reeb diameter $\mathsf{diam}_{\mathcal{R}}(\beta)$, the volume $\int_{X}\beta\wedge(d\beta)^{n}$, and the volume $vol_{(d\beta)^{n}}(\mathcal{T}(v_{\beta}))$ of the trajectory space. $\diamondsuit$
@@
$\int_{\gamma}(\beta+d\eta)=\int_{\gamma}\beta+(\eta(b)-\eta(a))=\int_{\gamma}\beta$
@@
[…]
@@
Also,
$vol_{(d\beta)^{n}}(\mathcal{T}(v_{\beta}))=\int_{\partial_{1}^{+}X(v_{\beta})}(d\beta)^{n}=\int_{\partial_{1}^{+}X(v_{\beta+d\eta})}(d\beta)^{n}$
$=\int_{\partial_{1}^{+}X(v_{\beta+d\eta})}(d(\beta+d\eta))^{n}=vol_{(d(\beta+d\eta))^{n}}(\mathcal{T}(v_{\beta+d\eta}))$
@@
Let $Y$ be a smooth manifold and $v$ a non-vanishing vector field on it that admits a Lyapunov function. Consider a compact smooth connected manifold $X$ with boundary, $\dim X=\dim Y$, and a regular imbedding $\Psi:X\to\mathsf{int}(Y)$ such that $v$ is boundary generic relative the hypersurface $\Psi(\partial_{1}X)$. We introduce a natural number $c^{\bullet}(\Psi,v)$ as half of the maximum number of times a $v$-trajectory intersects $\Psi(\partial_{1}X)$. $\diamondsuit$
@@
Even if the connected source and the target of $\Psi$ both carry convex vector fields, the number $c^{\bullet}(\Psi,v)$ can be arbitrary big. Indeed, just rap a very narrow ($\epsilon\ll 1$) band $Z=I\times[-\epsilon,\epsilon]$ with rounded corners $k$ times around the cylinder $C=S^{1}\times[-1,1]$, where $I$ denotes the interval $[0,1]$ and the vector field $v$ is aligned with the generators of the cylinder. Next, consider the embedding
$\Psi:X=S^{1}\times I\times[-\epsilon,\epsilon]\;\hookrightarrow\;S^{1}\times S^{1}\times[-1,1]=Y,$
obtained by multiplying the wrapping map $\psi:Z\subset C$ with the identity map on the first multiplayer $S^{1}$. Then $c^{\bullet}(\Psi,v_{Y})=k$. $\diamondsuit$
Consider a traversing $v$-flow on a compact manifold $Y$ and a regular embedding $\Psi:X\to\mathsf{int}(Y)$, where $X$ is a compact manifold of the same dimension as $Y$. Assuming that $v$ is boundary generic with respect to the hypersurface $\Psi(\partial_{1}X)$, with the help of $(-v)$-flow, we can generate a shadow $\mathcal{X}^{\partial}(\Psi)$ of $\Psi(X)$ on the boundary of $Y$. It is produced by forming the space $\mathcal{X}(\Psi)$ of $v$-trajectories through the points of $\Psi(X)$ (equivalently, through the locus $\partial_{1}^{+}\Psi(X)(v)$) and then taking the intersection $\mathcal{X}(\Psi)\cap\partial_{1}^{+}Y(v)$. This constructions presumes that $\Psi(X)$ is “semi-transparent” for the flow (see Fig. 1, the left diagram).
The next theorem resembles Gromov’s non-squeezing theorem [Gr2] in symplectic geometry, being transferred to the contact geometry environment (see Fig. 1). It deals with the embeddings $\Psi:X_{1}\to X_{2}$ ($\dim X_{1}=2n+1=\dim X_{2}$) that respect the contact forms: i.e., $\beta_{1}=\Psi^{\ast}(\beta_{2})$. It employs the number $c^{\bullet}(\Psi,v_{\beta_{2}})$ that measures how folded is the image $\Psi(\partial_{1}X_{1})$ with respest to the $v_{\beta_{2}}$-flow. The “equatorial” inequality of Theorem 6.2 may be also viewed as delivering a lower bound of the number $c^{\bullet}(\Psi,v_{\beta_{2}})$ in terms of the appropriate $\beta_{1}$- and $\beta_{2}$-generated symplectic $2n$-volumes.

### v6

@@ 5. On contact forms and contact vector fields on manifolds with boundary
A smooth vector field $u$ on $X$ is called contact field for a given contact $1$-form $\beta$, if $\mathcal{L}_{u}\beta=\lambda\cdot\beta$ for some smooth function $\lambda:X\to\mathbb{R}$. $\diamondsuit$
@@
To derive the next Theorem 5.1 we will rely on a couple of lemmas.
@@
Using that $\mathcal{L}_{v_{\beta}}\beta\equiv 0$ and $v_{\beta}\,\rfloor\,\beta=1$, this is equivalent to the condition $1\cdot dh+\mathcal{L}_{w}\beta=\lambda\cdot\beta,$ or, by Cartan’s formula and using that $w\,\rfloor\,\beta=0$, to the condition
@@
For $h$ such that $h^{-1}(0)=\partial_{1}X$, by Lemma 5.3, the vector field $u=h\cdot v_{\beta}+w$ is tangent to $\partial_{1}X$, since $h|_{\partial_{1}X}\equiv 0$. Thus, we getting the following claim.
@@
$\bullet$ Let $X=\{x^{2}+y^{2}+z^{2}\leq 1\}\subset\mathbb{R}^{3}$ and $\beta=dz+x\,dy$. The Reeb vector field $v_{\beta}$ is $\partial_{z}$. It is tangent to the $2$-sphere $\partial_{1}X$ along its equator $E=\{z=0\}\cap X$. Thus, the restriction of $\beta$ to $E$ is equal to the restriction of the form xdyx\,dy to $E$. If $\theta$ is the natural parameter on the circle $E$, then $\beta|_{E}=cos^{2}(\theta)\,d\theta$. It is non-negative everywhere on $E$ and vanishes at $\theta=\pm\pi/2$, or at $(x,y,z)=(0,\pm 1,0)$.
@@
The $u$-flow preserves the distribution $\xi_{\beta}$ and the sphere $\partial_{1}X$. It has exactly two fixed points, $(0,0,\pm 1)$.
$\bullet$ We pick the shell $\{1\leq x^{2}+y^{2}+z^{2}\leq 4\}$ for the role of $X$ and the same contact form $\beta=dz+x\,dy$. Then $\partial_{2}^{+}X(v_{\beta})$ is the equator $E_{1}=\{x^{2}+y^{2}+z^{2}=1\}\cap\{z=0\}$ and $\partial_{2}^{-}X(v_{\beta})$ is the equator $E_{2}=\{x^{2}+y^{2}+z^{2}=4\}\cap\{z=0\}$. The same computation leads to $\beta|_{E_{1}}\geq 0$ and $\beta|_{E_{2}}\leq 0$. In other words, $\pm\beta|_{\partial_{2}^{\pm}X(v_{\beta})}\geq 0$.
Now we consider the function $h=(x^{2}+y^{2}+z^{2}-4)(x^{2}+y^{2}+z^{2}-1)$. We leave to the reader to solve the equation $w\,\rfloor\,d\beta=-dh+dh(v_{\beta})\cdot\beta$ for $w\in\xi_{\beta}$.
@@
Therefore, $(\beta\wedge d\beta)|_{\partial_{2}^{-}X(v_{\beta})}\geq\;0.$ Note that $(\beta\wedge d\beta)||_{\partial_{2}^{-}X(v_{\beta})}=0$ along the circle $\{x_{1}=0\}\cap\{x_{2}=0\}\cap E$ in the $(y_{1},y_{2})$-plane. $\diamondsuit$
@@
  • For an even $j$,  $\beta|_{\mathsf{int}(\partial_{j}^{+}X(v_{\beta}))}$ is a contact form away from a set of measure zero. Moreover,
@@
  • For an odd $j$, $d\beta|_{\mathsf{int}(\partial_{j}^{+}X(v_{\beta}))}$ is a symplectic form away from a set of measure zero. Moreover,
    $\pm(d\beta)^{\frac{2n+1-j}{2}}\big|_{\partial_{j}^{\pm}X(v_{\beta})}\geq 0,\text{ and }(d\beta)^{\frac{2n+1-j}{2}}\big|\big|_{\partial_{j+1}X(v_{\beta})}\equiv 0\;.$
@@
Recall that the Gray Stability Theorem [Gray] claims that, for any smooth family $\{\xi_{t}\}_{t\in[0,1]}$ of contact structures on a closed smooth manifold $X$, there exists a smooth isotopy $\{\phi_{t}\}_{t\in[0,1]}$ such that $(\phi_{t})^{\ast}(\xi_{t})=\xi_{0}$ for all $t\in[0,1]$.
Lemma 5.5 (whose proof we skip) states a relative version of Gray’s theorem [Gray]. Its proof may be found in [EM], page 95.
@@
In fact, we can generalize a bit Lemma 5.5 by assuming that $\dot{\beta}_{t}||_{\partial_{1}X}=dh_{t}||_{\partial_{1}X}$, where $\dot{\beta}_{t}=_{\mathsf{def}}\frac{d}{dt}\beta_{t}$ and the functions $h_{t}$ are as in (4).
Lemma 5.6.
Let $X$ be a compact connected smooth $(2n+1)$-dimensional manifold $X$ with boundary. Let $\{h_{t}:\hat{X}\to\mathbb{R}\}_{t\in[0,1]}$ be a family of smooth functions, each of which has 0 as its regular value, $h_{t}^{-1}(0)=\partial_{1}X$, and $h_{t}^{-1}((-\infty,0])=X$.
Consider a family $\{\xi_{t}\}_{t\in[0,1]}$ of cooriented contact structures on $X$, generated by a smooth family of contact forms $\{\beta_{t}\}_{t\in[0,1]}$, such that the $t$-derivative $\dot{\beta}_{t}=dh_{t}$ along $\partial_{1}X$. [^5] The property $\dot{\beta}_{t}||_{\partial_{1}X}=dh_{t}||_{\partial_{1}X}$ implies that $\beta_{t}|_{\partial_{1}X}$ is constant.
Then there is a smooth $t$-family of diffeomorphism $\{\phi_{t}:X\to X\}_{t\in[0,1]}$, such that $(\phi_{t})^{\ast}(\xi_{t})=\xi_{0}$; in particular, $\phi_{1}$ is a contactomorphism that takes $\xi_{1}$ to $\xi_{0}$.
@@ 6. Non-squeezing of $2n$-symplectic volumes, Morse wrinkles of Reeb fields, and holography
Let $v_{\beta}$ be a traversing Reeb vector field for a contact $1$-form $\beta$ on $X$.
The Reeb diameter of $(X,\beta)$ is defined by
@@
where $\gamma$ runs over all $v_{\beta}$-trajectories. $\diamondsuit$
@@
Consider an orientation-preserving submersion $\Psi:X\to Y$ of two compact equidimentional manifolds. Let $\beta_{Y}$ be a contact form on $Y$ whose Reeb vector field $v_{Y}$ is traversing. We assume that the hypersurface $\Psi(\partial_{1}X)$ is boundary generic relative to $v_{Y}$.
@@
If $\Psi$ is a covering map (i.e., $\Psi$ is an immersion that is surjective), then
@@
Since $\Psi$ is submersion and $\Psi(\partial_{1}X)$ is boundary generic relative to $v_{Y}$, for any $v_{Y}$-trajectory $\gamma$ (which is a closed interval or a singleton) the set $\Psi^{-1}(\gamma)$ is either a disjoint union of finitely many closed intervals $\tilde{\gamma}$, some of which may be singletons. Moreover, $\Psi:\tilde{\gamma}\hookrightarrow\gamma$ is an immersion. Thus, $\int_{\tilde{\gamma}}\beta_{X}=\int_{\tilde{\gamma}}\Psi^{\ast}(\beta_{Y})=\int_{\Psi(\tilde{\gamma})}\beta_{Y}\leq\int_{\gamma}\beta_{Y}$. In the case of a covering $\Psi$, we get that $\Psi:\tilde{\gamma}\to\gamma$ is a diffeomorphism. In such a case, $\int_{\tilde{\gamma}}\beta_{X}=\int_{\gamma}\beta_{Y}$. Since any $v_{X}$-trajectory $\tilde{\gamma}$ is obtained from some $v_{Y}$-trajectory $\gamma$ as described above, the claims about the Reeb diameters follow. ∎
@@
The argument is a variation on the theme of Santalo formula [S]. For any $y\in\partial_{1}^{+}X(v_{\beta})$ we consider the $v$-trajectory $\gamma_{y}$ through $y$. The (v)(-v)-directed projection $p:X\to\partial_{1}^{+}X(v_{\beta})$ is a submersion, away from the set $p^{-1}(\partial_{2}X(v_{\beta}))$ of zero measure in $\partial_{1}^{+}X(v_{\beta})$. Thus, we can use the Dieudonne generalization of the Fubini formula [D]:
@@
together with Definition 6.1, to validate (6.2).
By Stokes’ theorem, applied to $\partial_{1}^{+}X(v_{\beta})$, (6.3) follows. Note that formula (6.3) testifies that $\int_{\partial_{2}X(v_{\beta})}\beta\wedge(d\beta)^{n-1}>0$, which supports weakly Conjecture 5.1.
Although these arguments are simple, it is somewhat surprising to realize the $(d\beta)^{n}$-volume of the trajectory space $\mathcal{T}(v_{\beta})$ depends only on the form $\beta\wedge(\partial\beta)^{n-1}|_{\partial_{2}X(v_{\beta})}$ on the “equator” $\partial_{2}X(v_{\beta})$. ∎
@@
By (6.3), $\mathsf{av}_{\mathcal{R}}(\beta)\;\leq\;\mathsf{diam}_{\mathcal{R}}(\beta).$ $\diamondsuit$
Example 7.1. Let $\beta=dz+x\,dy$ and let $X=\big\{\frac{x^{2}}{(A/2)^{2}}+\frac{y^{2}}{(B/2)^{2}}+\frac{z^{2}}{(C/2)^{2}}\leq 1\big\}$. Then the inequality (6.2) can be written as
@@
For any function $\eta\in C^{\infty}(\mathcal{T}(v_{\beta}))$, the contact form $\beta_{1}=\beta+d\eta$ shares with $\beta$ the Reeb diameter $\mathsf{diam}_{\mathcal{R}}(\beta)$, the volume $\int_{X}\beta\wedge(d\beta)^{n}$, and the volume $vol_{(d\beta)^{n}}(\mathcal{T}(v_{\beta}))$ of the trajectory space. $\diamondsuit$
@@
$\int_{\gamma}(\beta+d\eta)=\eta(b)-\eta(a)+\int_{\gamma}\beta=\int_{\gamma}\beta$
@@
$\int_{y\in\partial_{1}^{+}X(v_{\beta})}\Big(\int_{\gamma_{y}}\beta\Big)(d\beta)^{n}+\int_{y\in\partial_{1}^{+}X(v_{\beta})}\Big(\int_{\gamma_{y}}d\eta\Big)(d\beta)^{n}.$
Thanks to the property $\mathcal{L}_{v_{\beta}}\eta=0$, $\int_{\gamma_{y}}d\eta=0$ for all $y$, and the latter formula reduces to
@@
Also, using Lemma 4.7, we get
$vol_{(d\beta)^{n}}(\mathcal{T}(v_{\beta}))=_{\mathsf{def}}\int_{\partial_{1}^{+}X(v_{\beta})}(d\beta)^{n}=\int_{\partial_{1}^{+}X(v_{\beta+d\eta})}(d\beta)^{n}$
$\int_{\partial_{1}^{+}X(v_{\beta+d\eta})}(d(\beta+d\eta))^{n}=_{\mathsf{def}}vol_{(d(\beta+d\eta))^{n}}(\mathcal{T}(v_{\beta+d\eta}))$
@@
Let $Y$ be a smooth manifold and let $v$ a non-vanishing vector field on it that admits a Lyapunov function. Consider a compact smooth connected manifold $X$ with boundary, $\dim X=\dim Y$, and a regular imbedding $\Psi:X\to\mathsf{int}(Y)$ such that $v$ is boundary generic relative the hypersurface $\Psi(\partial_{1}X)$. We introduce a natural number $c^{\bullet}(\Psi,v)$ as one half of the maximal number of times a $v$-trajectory intersects $\Psi(\partial_{1}X)$. $\diamondsuit$
@@
Even if the connected source and the target of $\Psi$ both carry convex vector fields, the number $c^{\bullet}(\Psi,v)$ can be arbitrary big. Indeed, just rap a very narrow ($\epsilon\ll 1$) band $Z=I\times[-\epsilon,\epsilon]$ with rounded corners $k$ times around the cylinder $C=S^{1}\times[-1,1]$, where $I$ denotes the interval $[0,1]$ and the vector field $v$ is aligned with the generators of the cylinder. Next, consider the embedding Ψ: X = S^1 ×I ×[-ϵ, ϵ]  ↪ S^1 ×S^1 ×[-1, 1] =Y, obtained by multiplying the wrapping map $\psi:Z\subset C$ with the identity map on the first multiplayer $S^{1}$. Then $c^{\bullet}(\Psi,v_{Y})=k$. $\diamondsuit$
Consider a traversing $v$-flow on a compact manifold $Y$ and a regular embedding $\Psi:X\to\mathsf{int}(Y)$, where $X$ is a compact manifold of the same dimension as $Y$. Assuming that $v$ is boundary generic with respect to the hypersurface $\Psi(\partial_{1}X)$, with the help of (v)(-v)-flow, we can generate a shadow $\mathcal{X}^{\partial}(\Psi)$ of $\Psi(X)$ on the boundary of $Y$. It is produced by forming the space $\mathcal{X}(\Psi)$ of $v$-trajectories through the points of $\Psi(X)$ (equivalently, through the locus $\partial_{1}^{+}\Psi(X)(v)$) and then taking the intersection $\mathcal{X}(\Psi)\cap\partial_{1}^{+}Y(v)$. This constructions presumes that $\Psi(X)$ is “semi-transparent” for the flow (see Fig. 1, the left diagram).
The next theorem resembles Gromov’s non-squeezing theorem [Gr2] in symplectic geometry, being transferred to the contact geometry environment (see Fig. 1). It deals with the embeddings $\Psi:X_{1}\to X_{2}$ ($\dim X_{1}=2n+1=\dim X_{2}$) that respect the contact forms: i.e., $\beta_{1}=\Psi^{\ast}(\beta_{2})$. It employs the natural number $c^{\bullet}(\Psi,v_{\beta_{2}})$ that measures how folded is the image $\Psi(\partial_{1}X_{1})$ with respest to the $v_{\beta_{2}}$-flow. The “equatorial” inequality of Theorem 6.2 may be also viewed as delivering a lower bound of the number $c^{\bullet}(\Psi,v_{\beta_{2}})$ in terms of the appropriate $\beta_{1}$- and $\beta_{2}$-generated symplectic $2n$-volumes.